OpenStudy (alibaby):

What is the volume of a sphere with a radius of 9 inches? 108Ï€ cubic inches 242Ï€ cubic inches 642Ï€ cubic inches 972Ï€ cubic inches

OpenStudy (alibaby):

@zepdrix @Zarkon @Zackery17 @Zaynwich @HelpWithK12

OpenStudy (zackery17):

V≈3053.63in³

OpenStudy (alibaby):

huh

OpenStudy (alibaby):

none of the answer choices

KjSaif (kjsaif):

V = 4/3 π r^3 V = 4/3 π (9 inches) ^3 V = 972 π in^3